please help.
A line passing through (-10,-1)and (-8,y) is perpendicular to a line with slope -2/9. Find the value of y.
thank you.

Answer by scott8148(6628)

perpendicular lines have slopes that are negative reciprocals
this means that the slope of the line you are trying to find is 9/2

so, (y-(-1))/(-8-(-10))=9/2 ... (y+1)/2=9/2 ... y+1=9 ... y=8